# Search for Long-lived Particles at Future Lepton Colliders Using Deep Learning Techniques

### v5

@@ Abstract
Long-lived particles (LLPs) provide an unambiguous signal for physics beyond the Standard Model (BSM). They have a distinct detector signature, with decay lengths corresponding to lifetimes of around nanoseconds or longer. Lepton colliders allow LLP searches to be conducted in a clean environment, and such searches can reach their full physics potential when combined with machine learning (ML) techniques. This experimental study, utilizing comprehensive full simulation data samples, focuses on LLP searches resulting from Higgs decay in $e^{+}e^{-}\to ZH$. We demonstrate that, by employing deep neural network approaches the LLP signal efficiency can be improved up to 95% for an LLP mass around 50 GeV and a lifetime of approximately 1 nanosecond, while rejecting all SM backgrounds. Furthermore, the signal sensitivity for the branching ratio of Higgs decaying into LLPs reaches a state-of-art limit of $9.7\times 10^{-7}$ with a statistics of $4\times 10^{6}$ Higgs.
@@ 1 Introduction
In 2012, the discovery of the Higgs boson completed the final piece of the Standard Model (SM) [1, 2]. Remarkably, SM predictions align with almost all experimental observations. However, the SM does not address several critical questions in the universe’s evolution, such as the existence of dark matter [3, 4, 5], the matter-antimatter imbalance puzzle [6, 7] and the origin of the neutrino mass [8, 9]. Consequently, the search for new physics beyond the Standard Model (BSM) is both intriguing and necessary. Despite numerous efforts, BSM signals have remained elusive, suggesting that new physics may either require a higher energy threshold than currently achievable or that the coupling strength between SM and BSM particles is too weak to produce a statistically significant number of observable signals. Amidst this scenario, the hypothesis that BSM particles could possess long lifetimes, evading detection, has transitioned from a nascent idea to a widely accepted and vigorously pursued avenue within the physics community. These particles, often called long-lived particles (LLPs), serve as sensitive probes into BSM physics, such as supersymmetry (SUSY) theory [10, 11] or dark matter models [12]. LLPs could also provide insights for the matter-antimatter asymmetry [13] and the neutrino mass hierarchy, with their lifetimes varying widely. Recently, the search for LLPs has intensified across several experiments, such as ATLAS [14, 15], CMS [16, 17], and LHCb [18], as well as Belle II [19], BESIII [20], Babar [21], marking a keen interest in their potential to reveal new aspects of physics, even though definitive discoveries are yet to be made.
@@
Searching for LLPs using machine-learning (ML) assisted approaches at hadron colliders [22, 23, 24] has emerged as a promising strategy to surmount these experimental challenges. Such research employs ML techniques for object-level tagging to identify LLP signatures following comprehensive event reconstruction. The efficacy of this tagging is dependent on a constrained set of high-level input variables. The demanding environment at hadron colliders complicates background control, possibly necessitating neural networks with higher granularity and more extensive input features. Furthermore, the deployment of advanced triggers powered by rapid neural networks is crucial for the efficient detection of LLP-like events at hadron colliders [25, 26].
@@
Figure 1: Feynman diagrams of LLP production and decay. Two Feynman diagrams are presented illustrating the generation of long-lived particles (LLPs), denoted as $X$, through the Higgsstrahlung mechanism. On the left, the diagram shows the production of $XX$ followed by their subsequent decay into a $\nu\bar{\nu}$ pair and a $q\bar{q}$ pair, respectively, resulting in two jets. On the right, both $X$ decay into $q\bar{q}$ pairs, leading to the four jets final state.
@@
Electron-positron colliders, such as the Circular Electron Positron Collider (CEPC) [27] and the International Linear Collider (ILC) [28], provide clean collision environments with precise initial-state conditions. The Higgs boson production via $e^{+}e^{-}\rightarrow ZH$ at these facilities offers a clean channel to explore rare LLP decays, benefiting from well-defined initial kinematics and significantly reduced backgrounds compared to hadron colliders like the LHC [29, 30]. The primary goal of this analysis is to measure the cross section of LLP production through Higgs boson decays at future lepton colliders. Given that the SM Higgs production cross section at electron-positron colliders is accurately calculable and well-known, the LLP production cross section can be directly derived by multiplying the predicted SM $ZH$ cross section by the assumed branching ratios (BRs) of Higgs into LLPs. Therefore, these branching ratios are the fundamental parameters that we aim to constrain in this analysis.
@@
The Higgs boson can decay into a pair of long-lived particles (LLP, denoted as $X$), each of which subsequently decays into final state objects such as jets, charged leptons, or neutrinos. In the scope of our study, the search strategy is inspired by the glueball and Gauge Mediated Supersymmetry Breaking (GMSB) models [31, 32], where such LLPs naturally emerge and have accessible decay modes as defined by the process of $H\rightarrow XX\rightarrow\textrm{jets}$. This focus stems from the distinguishable detector signatures that the jet decay products can offer.
@@
Figure 1 shows Feynman diagrams for the LLP production process from the Higgs decay in lepton colliders with two jets (type I signal) or four jets (type II signal) in the final state. We focus on scenarios where the mixing between the new scalar particle $X$ and the Higgs boson is negligible. This assumption allows us to isolate and analyze specific decay channels of $X$ without the complexities introduced by universal couplings to all SM particles that would result from significant $X$-Higgs mixing. While $X$ could, in principle, decay into gauge bosons, our study focuses on the hadronic and invisible decay channels. This focus is motivated by scenarios such as effective $Z^{\prime}$ models, where the new scalar $X$ couples preferentially to quarks and neutrinos over charged leptons and gauge bosons [33, 34, 35]. The invisible decay channel of $X$ is defined to include any final states that result in missing energy signatures, such as neutrinos or potential dark matter candidates. This approach aligns with various BSM scenarios where $X$ serves as a mediator between the SM and the dark sector, leading to missing energy signatures in the detector. Additionally, hadronic and invisible final states pose significant experimental challenges due to large QCD backgrounds and missing energy signatures, respectively, making them compelling channels for detailed study.
@@
Other potential decay channels such as $H\rightarrow XX\rightarrow\textrm{invisible}$ and $H\rightarrow XX\rightarrow\textrm{leptons}$ are not considered in this study. Particularly, the $H\rightarrow XX\rightarrow\textrm{invisible}$ channel, classified as a type III signal, results in decay products that leave no detectable signature within the detector’s sensitivity range. Such invisibly decaying LLPs do not contribute to observable events that can be analyzed within our current experimental framework. Consequently, these decays are more appropriately addressed in analyses specifically dedicated to $H\rightarrow\textrm{invisible}$ decays [36]. Similarly, while $H\rightarrow XX\rightarrow\textrm{leptons}$ presents an interesting decay channel, the current study prioritizes jet signatures due to their robust detection capabilities and the analytical focus on hadronic final states. Future research may expand to encompass these additional decay modes, taking into account different event topologies and analysis methodologies.
@@
In this study, we investigate the feasibility and sensitivity of employing advanced machine learning methods, particularly deep learning techniques such as Convolutional Neural Networks (CNNs) and Graph Neural Networks (GNNs), to learn the intricate topologies and kinematic features of LLPs. We conduct a comparative analysis of the two distinct machine learning architectures to evaluate their respective performances in reconstructing LLP signals. Each model is independently trained and assessed, providing insights into their individual capabilities and limitations within the context of LLP detection. By applying these methods directly to low-level detector data, we achieve remarkable improvements in LLPs signal reconstruction efficiency and background suppression. The search sensitivity for LLPs at future $e^{+}e^{-}$ colliders achieved using deep learning techniques significantly surpasses that of traditional methods, establishing state-of-the-art limits on the Higgs branching ratio into LLPs. This approach enables us to derive stringent constraints on the branching ratios, thus enhancing our ability to identify potential LLP signals and probe BSM physics effectively. Importantly, the ML-based methodology we present offers broad applicability, promising an easy adaptation for future lepton collider experiments, including the CEPC [27], ILC [28], FCC [37], and CLIC [38].
@@ 2 Signal and Background Samples
In this study, we utilize fully simulated Monte Carlo (MC) samples for both the signal and background analyses. To generate the long-lived particle (LLP) signal samples, MadGraph 3.0.1 [39] is employed to cover both type I and type II signals across various mass points (50 GeV, 10 GeV, and 1 GeV) and lifetimes (0.001 ns, 0.1 ns, 1 ns, 10 ns, 100 ns). A total of $1.5\times 10^{7}$ events are produced, distributing $10^{6}$ events for each combination of mass point and lifetime.
The backgrounds predominantly consist of SM processes resulting in jets in the final state, notably $e^{+}e^{-}\rightarrow\bar{q}q$ and $e^{+}e^{-}\rightarrow VV$, where $VV$ includes both $WW$ and $ZZ$. Additionally, the SM process $e^{+}e^{-}\rightarrow ZH\rightarrow\textrm{inclusive}$ is also considered, given its significance as the main production mode for SM Higgs. These SM processes are simulated using Whizard 1.95 [40, 41]. A total of $10^{7}$ events is simulated for both the $e^{+}e^{-}\rightarrow\bar{q}q$ and $e^{+}e^{-}\rightarrow VV$ processes, while the $e^{+}e^{-}\rightarrow ZH\rightarrow\textrm{inclusive}$ process has $10^{6}$ events simulated. The top quark and tri-boson related backgrounds have negligible cross sections at a center-of-mass energy of 240 GeV and therefore are not considered. Minor backgrounds such as pileup background and cosmic ray background are also neglected in this analysis (Appendix B). Details on the cross-sections and simulated statistics for the LLP signal and SM background are provided in Table 1.
For each Monte Carlo (MC) sample, a full detector simulation is performed following the CEPC conceptual design [42]. The CEPC detector comprises a silicon-based vertexing and tracking system, a Time Projection Chamber (TPC) tracker, a high-granularity calorimetry system, a 3 Tesla superconducting solenoid magnet, and a muon detection system integrated within the magnet’s iron return yoke. The calorimetry system consists of an electromagnetic calorimeter (ECAL) and an iron-scintillator hadronic calorimeter (HCAL). Simulations are carried out using MOKKAC software [43], a dedicated Geant4-based tool customized for the CEPC framework, which accounts for comprehensive particle showering effects and calibrated detector responses. The calorimeter resolutions are modeled to achieve an energy resolution of $16\%/\sqrt{E/\textrm{GeV}}\oplus 1\%$ for the ECAL and $60\%/\sqrt{E/\textrm{GeV}}\oplus 1\%$ for the HCAL and muon detector. In addition, detector hits are assigned a time resolution of 1 nanosecond. Following full detector simulation, no further event or object reconstruction is applied; rather, the digitized detector hits are directly utilized as inputs for the subsequent ML-based analyses.
@@
Process LLP signal SM ZH $q\bar{q}$ $ZZ$ $WW$
$\sigma[\mathrm{fb}]$ 203.66 54106.86 1110.37 16721.77
Statistics $1.5\times 10^{7}$ $1.0\times 10^{6}$ $1.0\times 10^{7}$ $1.1\times 10^{6}$ $8.9\times 10^{6}$
@@ 3.1 General Analysis Strategy
Initially, a signal acceptance selection is applied to ensure that at least one LLP is within the geometric boundaries of the CEPC detector [42], with acceptance factors detailed in Table 2. Subsequently, signal events are classified based on the number of LLP decays detectable within the detector and their consequent final states. This classification yields three categories: $\mathcal{D}_{0}$, $\mathcal{D}_{1}$, and $\mathcal{D}_{2}$, denoting events with 0, 1, and 2 detectable LLPs leading to jet final states, respectively.
@@
The classification output is a set of prediction scores, indicating the probability of an event belonging to each of the five categories. Following this NN-based classification, an XGBoost algorithm [44] is implemented to use these prediction scores as inputs to generate a single output score. This score discriminates between signal events and SM background.
@@ 3.2 Convolutional Neural Network
For the CNN approach, each event with low-level detector information is transformed into a two-channel image with $200\times 200$ pixels in $(R,\phi)$, where $R$ represents the distance of the position of the hits to the interaction point ($0-6$ m) and $\phi$ represents the azimuthal angle of the hits ($0-2\pi$). The first channel (energy channel) represents the sum of all hits energies in a pixel. The second channel (time channel) represents the time difference ($\Delta t$) associated with the most energetic hit in a pixel as shown in Eq. 1.
@@
$t_{hit,maxE}$ is the time of the hit with the maximum energy in a pixel, $r_{hit,maxE}$ is the Euclidean distance from the interaction point to the hit location, and $c$ is the speed of light in vacuum. A selection of hits energy larger than $0.1$ GeV is placed to suppress contributions from extremely low energy hits of secondary particles. Figure 3 shows images of hits for signal and background events. The event images illustrate the discriminating power of CNN to separate LLP signals with displaced energies or objects from SM backgrounds. By feeding event images to ResNet18 neural networks [45] shown in Figure 4, a multi-label classification is performed. The CNN model is trained using 5 GPUs with a batch size of 256 per GPU. The learning rate is adaptively set, beginning at $10^{-4}$ and progressively decreasing to $10^{-6.5}$. The training employs the Adam optimizer with a weight decay of 0.01. The training duration is set at 10 epochs, and the model’s final configuration is determined by the epoch that achieves the lowest validation loss. The distributions of the training and testing losses for CNN are shown in Appendix C.

### v6

@@ Abstract
Long-lived particles (LLPs) provide an unambiguous signal for physics beyond the Standard Model (BSM). They have a distinct detector signature, with decay lengths corresponding to lifetimes of around nanoseconds or longer. Lepton colliders allow LLP searches to be conducted in a clean environment, and such searches can reach their full physics potential when combined with machine learning (ML) techniques. This experimental study, utilizing comprehensive full simulation data samples, focuses on LLP searches resulting from Higgs decay in $e^{+}e^{-}\to ZH$. We demonstrate that, by employing deep neural network approaches the LLP signal efficiency can be improved up to 95% for an LLP mass around 50 GeV and a lifetime of approximately 1 nanosecond, while rejecting all SM backgrounds. Furthermore, the signal sensitivity for the branching ratio of Higgs decaying into LLPs reaches a state-of-art limit of $1.0\times 10^{-6}$ with a statistics of $4\times 10^{6}$ Higgs.
@@ 1 Introduction
In 2012, the discovery of the Higgs boson completed the final piece of the Standard Model (SM) [CMS:2012qbp, CMS:2013btf]. Remarkably, SM predictions align with almost all experimental observations. However, the SM does not address several critical questions in the universe’s evolution, such as the existence of dark matter [Kaplan:2009ag, Dienes:2011ja, Kim:2013ivd], the matter-antimatter imbalance puzzle [Bouquet:1986mq, Campbell:1990fa] and the origin of the neutrino mass [Graesser:2007yj, Maiezza2015]. Consequently, the search for new physics beyond the Standard Model (BSM) is both intriguing and necessary. Despite numerous efforts, BSM signals have remained elusive, suggesting that new physics may either require a higher energy threshold than currently achievable or that the coupling strength between SM and BSM particles is too weak to produce a statistically significant number of observable signals. Amidst this scenario, the hypothesis that BSM particles could possess long lifetimes, evading detection, has transitioned from a nascent idea to a widely accepted and vigorously pursued avenue within the physics community. These particles, often called long-lived particles (LLPs), serve as sensitive probes into BSM physics, such as supersymmetry (SUSY) theory [Arvanitaki:2012ps, Fan:2011yu] or dark matter models [Bertone:2016nfn]. LLPs could also provide insights for the matter-antimatter asymmetry [Dine:2003ax] and the neutrino mass hierarchy, with their lifetimes varying widely. Recently, the search for LLPs has intensified across several experiments, such as ATLAS [ATLAS_llp_muon, ATLAS2024LLP], CMS [CMS:llp2023, CMS_llp_2024], and LHCb [Calefice:2022upw], as well as Belle II [Ferber:2022rsf], BESIII [2023137698], Babar [BaBar2015], marking a keen interest in their potential to reveal new aspects of physics, even though definitive discoveries are yet to be made.
@@
Searching for LLPs using machine-learning (ML) assisted approaches at hadron colliders [Cui:2014twa, Bhattacherjee:2019fpt, CMS:2024trg] has emerged as a promising strategy to surmount these experimental challenges. Such research employs ML techniques for object-level tagging to identify LLP signatures following comprehensive event reconstruction. The efficacy of this tagging is dependent on a constrained set of high-level input variables. The demanding environment at hadron colliders complicates background control, possibly necessitating neural networks with higher granularity and more extensive input features. Furthermore, the deployment of advanced triggers powered by rapid neural networks is crucial for the efficient detection of LLP-like events at hadron colliders [Alimena2020LLP, Coccaro2023LLP].
@@
Figure 1: Feynman diagrams of LLP production and decay. Two Feynman diagrams are presented illustrating the generation of LLPs, denoted as $X$, through the Higgsstrahlung mechanism. On the left, the diagram shows the production of XXXX followed by their subsequent decay into a $\nu\bar{\nu}$ pair and a $q\bar{q}$ pair, respectively, resulting in two jets. On the right, both $X$ decay into $q\bar{q}$ pairs, leading to the four jets final state.
@@
Electron-positron colliders, such as the Circular Electron Positron Collider (CEPC) [CEPCStudyGroup:CDR1] and the International Linear Collider (ILC) [ILC:2019gyn], provide clean collision environments with precise initial-state conditions. The Higgs boson production via $e^{+}e^{-}\rightarrow ZH$ at these facilities offers a clean channel to explore rare LLP decays, benefiting from well-defined initial kinematics and significantly reduced backgrounds compared to hadron colliders like the LHC [ATLAS:2018tup, CMS:2020iwv]. The primary goal of this analysis is to measure the cross section of LLP production through Higgs boson decays at future lepton colliders. Given that the SM Higgs production cross section at electron-positron colliders is accurately calculable and well-known, the LLP production cross section can be directly derived by multiplying the predicted SM $ZH$ cross section by the assumed branching ratios (BRs) of Higgs into LLPs. Therefore, these branching ratios are the fundamental parameters that we aim to constrain in this analysis.
@@
The Higgs boson can decay into a pair of long-lived particles (LLP, denoted as $X$), each of which subsequently decays into final state objects such as jets, charged leptons, or neutrinos. In the scope of our study, the search strategy is inspired by the glueball and Gauge Mediated Supersymmetry Breaking (GMSB) models [Giudice:1998bp, Meade:2010ji], where such LLPs naturally emerge and have accessible decay modes as defined by the process of $H\rightarrow XX\rightarrow\textrm{jets}$. This focus stems from the distinguishable detector signatures that the jet decay products can offer.
@@
Figure 1 shows Feynman diagrams for the LLP production process from the Higgs decay in lepton colliders with two jets (type I signal) or four jets (type II signal) in the final state. We focus on scenarios where the mixing between the new scalar particle $X$ and the Higgs boson is negligible. This assumption allows us to isolate and analyze specific decay channels of $X$ without the complexities introduced by universal couplings to all SM particles that would result from significant $X$-Higgs mixing. While $X$ could, in principle, decay into gauge bosons, our study focuses on the hadronic and invisible decay channels. This focus is motivated by scenarios such as effective $Z^{\prime}$ models, where the new scalar $X$ couples preferentially to quarks and neutrinos over charged leptons and gauge bosons [Fox:2011qd, Ilten:2016tkc, Ilten:2018crw]. The invisible decay channel of $X$ is defined to include any final states that result in missing energy signatures, such as neutrinos or potential dark matter candidates. This approach aligns with various BSM scenarios where $X$ serves as a mediator between the SM and the dark sector, leading to missing energy signatures in the detector. Additionally, hadronic and invisible final states pose significant experimental challenges due to large QCD backgrounds and missing energy signatures, respectively, making them compelling channels for detailed study.
@@
Other potential decay channels such as $H\rightarrow XX\rightarrow\textrm{invisible}$ and $H\rightarrow XX\rightarrow\textrm{leptons}$ are not considered in this study. Particularly, the $H\rightarrow XX\rightarrow\textrm{invisible}$ channel, classified as a type III signal, results in decay products that leave no detectable signature within the detector’s sensitivity range. Such invisibly decaying LLPs do not contribute to observable events that can be analyzed within our current experimental framework. Consequently, these decays are more appropriately addressed in analyses specifically dedicated to $H\rightarrow\textrm{invisible}$ decays [tan_2020]. Similarly, while $H\rightarrow XX\rightarrow\textrm{leptons}$ presents an interesting decay channel, the current study prioritizes jet signatures due to their robust detection capabilities and the analytical focus on hadronic final states. Future research may expand to encompass these additional decay modes, taking into account different event topologies and analysis methodologies.
@@
In this study, we investigate the feasibility and sensitivity of employing advanced machine learning methods, particularly deep learning techniques such as Convolutional Neural Networks (CNNs) and Graph Neural Networks (GNNs), to learn the intricate topologies and kinematic features of LLPs. We conduct a comparative analysis of the two distinct machine learning architectures to evaluate their respective performances in reconstructing LLP signals. Each model is independently trained and assessed, providing insights into their individual capabilities and limitations within the context of LLP detection. By applying these methods directly to low-level detector data, we achieve remarkable improvements in LLPs signal reconstruction efficiency and background suppression. The search sensitivity for LLPs at future $e^{+}e^{-}$ colliders achieved using deep learning techniques significantly surpasses that of traditional methods, establishing state-of-the-art limits on the Higgs branching ratio into LLPs. This approach enables us to derive stringent constraints on the branching ratios, thus enhancing our ability to identify potential LLP signals and probe BSM physics effectively. Importantly, the ML-based methodology we present offers broad applicability, promising an easy adaptation for future lepton collider experiments, including the CEPC [CEPCStudyGroup:CDR1], ILC [ILC:2019gyn], FCC [FCC:2018byv], and CLIC [CLIC:2018fvx].
@@ 2 Signal and Background Samples
In this study, we utilize fully simulated Monte Carlo (MC) samples for both the signal and background analyses. To generate the long-lived particle (LLP) signal samples, MadGraph 3.0.1 [Alwall:2014hca] is employed to cover both type I and type II signals across various mass points (50 GeV, 10 GeV, and 1 GeV) and lifetimes (0.001 ns, 0.1 ns, 1 ns, 10 ns, 100 ns). A total of $1.5\times 10^{7}$ events are produced, distributing $10^{6}$ events for each combination of mass point and lifetime.
The backgrounds predominantly consist of SM processes resulting in jets in the final state, notably $e^{+}e^{-}\rightarrow\bar{q}q$ and $e^{+}e^{-}\rightarrow VV$, where $VV$ includes both $WW$ and $ZZ$. Additionally, the SM process $e^{+}e^{-}\rightarrow ZH\rightarrow\textrm{inclusive}$ is also considered, given its significance as the main production mode for SM Higgs. These SM processes are simulated using Whizard 1.95 [Kilian:2007gr, Moretti:2001zz]. A total of $10^{7}$ events is simulated for both the $e^{+}e^{-}\rightarrow\bar{q}q$ and $e^{+}e^{-}\rightarrow VV$ processes, while the $e^{+}e^{-}\rightarrow ZH\rightarrow\textrm{inclusive}$ process has $10^{6}$ events simulated. The top quark and tri-boson related backgrounds have negligible cross sections at a center-of-mass energy of 240 GeV and therefore are not considered. Minor backgrounds such as pileup background and cosmic ray background are also neglected in this analysis (Appendix B). Details on the cross-sections and simulated statistics for the LLP signal and SM background are provided in Table 1.
For each MC sample, a full detector simulation is performed following the CEPC conceptual design [CEPCStudyGroup:CDR2]. The CEPC detector comprises a silicon-based vertexing and tracking system, a Time Projection Chamber (TPC) tracker, a high-granularity calorimetry system, a 3 Tesla superconducting solenoid magnet, and a muon detection system integrated within the magnet’s iron return yoke. The calorimetry system consists of an electromagnetic calorimeter (ECAL) and an iron-scintillator hadronic calorimeter (HCAL). Simulations are carried out using MOKKAC software [CEPC-SIMU-2017-001], a dedicated Geant4-based tool customized for the CEPC framework, which accounts for comprehensive particle showering effects and calibrated detector responses. The calorimeter resolutions are modeled to achieve an energy resolution of $16\%/\sqrt{E/\textrm{GeV}}\oplus 1\%$ for the ECAL and $60\%/\sqrt{E/\textrm{GeV}}\oplus 1\%$ for the HCAL and muon detector. In addition, detector hits are assigned a time resolution of 1 nanosecond. Following full detector simulation, no further event or object reconstruction is applied; rather, the digitized detector hits are directly utilized as inputs for the subsequent ML-based analyses.
@@
Process LLP signal SM ZH $q\bar{q}$ ZZZZ WWWW
$\sigma[\mathrm{fb}]$ - 203.66 54106.86 1110.37 16721.77
Statistics $1.5\times 10^{7}$ $\times 10^{6} 1.0×10$$\times 10^{7}$ $1.1\times 10^{6}$ $8.9\times 10^{6}$
@@ 3.1 General Analysis Strategy
Initially, a signal acceptance selection is applied to ensure that at least one LLP is within the geometric boundaries of the CEPC detector [CEPCStudyGroup:CDR2], with acceptance factors detailed in Table 2. Subsequently, signal events are classified based on the number of LLP decays detectable within the detector and their consequent final states. This classification yields three categories: $\mathcal{D}_{0}$, $\mathcal{D}_{1}$, and $\mathcal{D}_{2}$, denoting events with 0, 1, and 2 detectable LLPs leading to jet final states, respectively.
@@
The classification output is a set of prediction scores, indicating the probability of an event belonging to each of the five categories. Following this NN-based classification, an XGBoost algorithm [xgboost] is implemented to use these prediction scores as inputs to generate a single output score. This score discriminates between signal events and SM background.
@@ 3.2 Convolutional Neural Network
For the CNN approach, each event with low-level detector information is transformed into a two-channel image with $200\times 200$ pixels in $(R,\phi)$, where $R$ represents the distance of the position of the hits to the interaction point (060-6 m) and $\phi$ represents the azimuthal angle of the hits ($0-2\pi$). The first channel (energy channel) represents the sum of all hits energies in a pixel. The second channel (time channel) represents the time difference ($\Delta t$) associated with the most energetic hit in a pixel as shown in Eq. 1.
@@
$t_{hit,maxE}$ is the time of the hit with the maximum energy in a pixel, $r_{hit,maxE}$ is the Euclidean distance from the interaction point to the hit location, and $c$ is the speed of light in vacuum. A selection of hits energy larger than $0.1$ GeV is placed to suppress contributions from extremely low energy hits of secondary particles. Figure 3 shows images of hits for signal and background events. The event images illustrate the discriminating power of CNN to separate LLP signals with displaced energies or objects from SM backgrounds. By feeding event images to ResNet18 neural networks [DBLP:journals/corr/HeZRS15] shown in Figure 4, a multi-label classification is performed. The CNN model is trained using 5 GPUs with a batch size of 256 per GPU. The learning rate is adaptively set, beginning at $10^{-4}$ and progressively decreasing to $10^{-6.5}$. The training employs the Adam optimizer with a weight decay of 0.01. The training duration is set at 10 epochs, and the model’s final configuration is determined by the epoch that achieves the lowest validation loss. The distributions of the training and testing losses for CNN are shown in Appendix C.